OpenStudy (anonymous):

Solve the equation for y.Then find the value of y for the given value of x. 3x-6y=6; x=2

OpenStudy (yamyam70):

Hi, First we transpose -6y to the right side, and 6 to the left side. Then we substitute the value of x.

OpenStudy (yamyam70):

welcome to open study :)

OpenStudy (yamyam70):

Here is the link for the code of conduct , http://openstudy.com/code-of-conduct :)

OpenStudy (akashdeepdeb):

Did you understand @alexaaa ? :)

OpenStudy (anonymous):

so how would it be set up? Since its 6-6y=6

OpenStudy (anonymous):

not really, just a little. @AkashdeepDeb

OpenStudy (yamyam70):

3x-6y=6; , you know that when we transfer "terms" we change signs right?

OpenStudy (akashdeepdeb):

3x-6y=6 Write the equation by interchanging LHS AND RHS 3x - 6 = 6y y = (3x-6)/6 y = 1x/2 - 1 Okay? So this is y in terms of x Now find y value by substituting x=2 in place of x! :D Understood? :)

OpenStudy (anonymous):

Yes! Thank you. You explained it better than my teacher.

OpenStudy (yamyam70):

Don't forget to close the question :)

OpenStudy (akashdeepdeb):

:)